OpenStudy (anonymous):

Change this radical to an algebraic expression with fractional exponents. (sqrt)b

OpenStudy (bibby):

You can rewrite roots as fractional exponents. For example \(\large \sqrt[3]{x+2}={(x+2)}^\frac{1}{3}\)

OpenStudy (anonymous):

I guess i still dont see how to do it with b

OpenStudy (lochana):

as bubby mention before , you can write (sqrt)b as \[b^{1/2}\]

OpenStudy (anonymous):

But how do you get to that i guess is my question.

OpenStudy (lochana):

hmm.. do you know what sqrt(100) is ? it's 10 . so we can write \[100^{1/2}\] = 10

OpenStudy (anonymous):

Okay.

OpenStudy (lochana):

got it?

OpenStudy (anonymous):

kinda